Whispering Company purchases equipment on January 1, Year 1, at a cost of $534,000. The asset is expected to have a service life of 12 years and a salvage value of $48,060 


Compute the amount of depreciation for each of Years 1 through 3 using the straight-line depreciation method. (Round answers to 0 decimal places, e.g. 5,125.) 

Depreciation for Year 1 $ 

Depreciation for Year 2 $ 

Depreciation for Year 3 $ 


Compute the amount of depreciation for each of Years 1 through 3 using the sum-of-the-years'-digits method. 

Depreciation for Year 1 $ 

Depreciation for Year 2 $

Depreciation for Year 3 $
Answer #1

1. Annual depreciation = (534000-48060)/12 = $40495

Depreciation for year 1 = $40495

Depreciation for year 2 = 40495

Depreciation for year 3 = 40495

2. Total sum = 1 to 12 = 78

Depreciation per digit = (534000-48060)/78 = $6230

Depreciation for year 1 = 6230*12 = 74760

Depreciation for year 2 = 6230*11 = 68530

Depreciation for year 3 =6230*10 = 62300

> Would you add the solution with double-declining balance method ?

JPate Mon, Jun 21, 2021 4:55 PM